1.use the two companies NKE and UAA.What are the respective betas for NKE and UAA. Use YahooFinance.com to determine this.

2) Assuming that the risk-free rate is 1.5% and the historical return on the S&P 500 is 11%, what is the required return of NKE & UAA’s equity? Use the CAPM.
Answer #1

NKE Beta = 0.83

UAA Beta = 0.53

Required rate of return as per CAPM = risk-free rate + (historical return on the S&P 500 - risk-free rate) * beta

Required rate of return:

NKE = 1.5 + (11-1.5) * 0.83 = 9.39%

UAA = 1.5 + (11-1.5) * 0.53 = 6.54%

NKE requires a higher rate of return compared to UAA due to NKE's higher beta.